Schedule of Activity Costs
Quality Control Activities Activity Cost
Process audits $49,500
Training of machine operators 27,400
Processing returned products 18,100
Scrap processing (disposal) 25,300
Rework 7,900
Preventative maintenance 27,100
Product design 41,100
Warranty work 9,100
Finished goods inspection 23,300

From the above schedule of activity costs, determine the appraisal costs.

Schedule of Activity Costs
Quality Control Activities Activity Cost
Process audits $49,200
Training of machine operators 28,200
Processing returned products 18,900
Scrap processing (disposal) 26,600
Rework 6,700
Preventative maintenance 32,000
Product design 40,800
Warranty work 7,200
Finished goods inspection 23,800

From the above schedule of activity costs, determine the total activity cost.
Answer #1

From the above schedule of activity costs, determine the appraisal costs.

Process audits + Processing returned products + Rework = 49500 + 18100 + 7900 = $75500

From the above schedule of activity costs, determine the total activity cost.

Process audits + Training of machine operators + Processing returned products + Processing returned products + Rework + Preventative maintenance + Product design + Warranty work + Finished goods inspection

49200 + 28200 + 18900 + 26600 + 6700 + 32000 + 40800 + 7200 + 23800 = $233400